# KCL-TH-2024-59

### v1

@@ VI Discussion
The most striking result is the fact that BS mergers at high compactness generate more gravitational radiation than equivalent mergers of BHs. This was first noted in studies of solitonic models in Refs. Helfer et al. (2022); Evstafyeva et al. (2023a) and in Ref. Helfer et al. (2019) for the $m^{2}\phi^{2}$ model of a real scalar field (sometimes called an oscillaton). The fact that it appears in both real and complex scalar models suggests that head-on mergers of compact bosonic self-gravitating stars generate significantly more GW energy than equivalent BH mergers. It is thus an interesting question to ask whether this result extends to inspiral mergers of BSs. Indeed, our results demonstrate that, above some model-dependent compactness, BS mergers are always louder than BHs, even if they eventually approach (but do not reach) BH compactness – “deformable BSs are louder than stiff BHs”. Our conjecture for this result – supported by the numerical observations of “needles” in the GW power during mergers of unstable Branch IIUM mergers – is that during the collision phase, the increased deformation in a less compact BS results in a more strongly perturbed and asymmetric merger remnant whose subsequent relaxation generates more GWs. It will be interesting to see if this result extends to other types of horizonless compact objects such as Proca stars or perhaps even NSs.
@@
In general, BS models become unstable at high (but below BH) compactness where perturbations can trigger a rapid collapse to a BH. In both mini-BS and oscillaton models, this represents the only delineation point between stable and unstable BSs. However, in more complicated models with two or more minima in the potential, the presence of additional stable and unstable equilibrium points in the potential $V(|\phi_{\mathrm{c}}|)$ can introduce additional structure in the stability diagram (see e.g. Ref. Helfer et al. (2017)). For the solitonic models we have studied, this manifests itself in the form of additional stable/unstable branches in the solution space of BSs (see Fig. 4). In these additional unstable regions, a BS may either collapse to a BH or migrate to a different BS. In all cases, the existence of these unstable regions implies that, in the limit of infinite infall time, the GW energy $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is discontinuous at the points of transition from stable to unstable branches. Unfortunately, since we do not measure GW energies over extraction spheres directly, but only the strain at one point on the sphere, there is scant hope for observing such discontinuities.
@@
In section IV, we have studied in detail the nature of the instability, in particular the time-scale or migration time of the transition. Furthermore, if the instability leads to a transition into another (excited) BS which oscillates, then we should expect “needles” structure in the energy function $E_{\rm GW}(|\phi_{\mathrm{c}}|)$. The main lesson here is that, there can be rapid changes in GW energy over small changes in the parameter space. For example, if we merge two equal mass oscillating BSs, the GW emitted would depend strongly on the phase of the pulsation at merger, complicating any observational inference over its progenitors. This raises two further important questions. (i) What is the timescale of relaxation of dynamic BS configurations to an equilibrium state? Numerical studies often result in long-lived dynamical stars (see e.g. Refs. Croft et al. (2023); Siemonsen and East (2023b) or our Fig. 5), but is it long enough to be of relevance for inspirals? (ii) What kind of environmental effects could induce dynamic perturbations of the kind experienced by our unstable migrating BSs?
@@
Finally, any program to search for BSs in the observational data stream requires the construction of waveform models or templates akin to the IMRPhenom, SEOBNR or NRSur approximants for BH-binary mergers. Our results suggest that the GWs from the BS star merger phase can exhibit significant differences in the GW amplitude as compared to BH mergers; cf. Fig. 9. The GW amplitude, however, is essentially degenerate with the luminosity distance and thus of little help for characterizing the nature of the source. This naturally raises the question whether the GW signature of the needle regime may exhibit more structural features in inspirals and/or for other BS parameters like unequal mass ratios or non-zero spins. Depending on the longevity of BS perturbations, comprehensive binary BS templates may need to account for these effects, perhaps using phenomenological parameters to extend simple semi-analytic quadrupole models (e.g. Ref. Toubiana et al. (2021)) with rigid mass distributions. Given the complexity of the merger physics, however, we suspect that surrogate models, as for example constructed for Proca-star head-on mergers in Ref Luna et al. (2024), may provide the most efficient way forward.
@@ Acknowledgements.
Appendix A Gravitational Wave Extraction
Numerical relativity simulations in $3+1$ dimensions commonly use the Newman-Penrose scalar $\Psi_{4}$ to extract gravitational radiation. Here we follow the discussion of GW extraction in $d+1$ dimensions of Ref. Cook and Sperhake (2017) and apply it to our case of $2+1$ evolutions ($d=2$). The outgoing gravitational radiation is encoded in the the projections of the Weyl tensor $C_{ABCD}$ onto a null tetrad $\{k^{A},l^{A},m^{A},\bar{m}^{A}\}$
@@ Appendix A Gravitational Wave Extraction
here capital Latin indices run from 00 to $d$. In order to construct the null tetrad, we first define an orthonormal basis $\{e_{(0)}^{A},e_{(1)}^{A},e_{(2)}^{A},e_{(3)}^{A}\}$, where $e_{(0)}^{A}$ is the unit normal to the spatial hypersurfaces, $e_{(1)}^{A}$ is the unit normal radial vector and $e_{(2)}^{A},e_{(3)}^{A}$ are the angular vectors, which are typically constructed via Gram-Schmidt orthonormalisation. We then pick the tetrad consisting of ingoing and outgoing null vectors, $k^{A}$ and $l^{A}$ respectively, and a complex null vector $m^{A}$ and its conjugate $\bar{m}^{A}$,
@@
Following the notation of Ref. Cook and Sperhake (2017), we will denote the off-domain components with a ’$w$’ index, so that in $2+1$ we have, $e_{(3)}^{A}\rightarrow e_{(w)}^{A}$. It is then straightforward to show that the outgoing gravitational radiation in Eq. (5) can be re-expressed by the real quantity
@@
where $\Omega_{22}$ and $\Omega_{ww}$ can be obtained directly from Eqs.(4.22)-(4.35) of Ref. Cook and Sperhake (2017).

### v2

@@ VI Discussion
The most striking result is the fact that BS mergers at high compactness generate more gravitational radiation than equivalent mergers of BHs. This was first noted in studies of solitonic models in Refs. Helfer:2021brt; Evstafyeva:2022bpr and in Ref. Helfer:2018vtq for the $m^{2}\phi^{2}$ model of a real scalar field (sometimes called an oscillaton). The fact that it appears in both real and complex scalar models suggests that head-on mergers of compact bosonic self-gravitating stars generate significantly more GW energy than equivalent BH mergers. It is thus an interesting question to ask whether this result extends to inspiral mergers of BSs. Indeed, our results demonstrate that, above some model-dependent compactness, BS mergers are always louder than BHs, even if they eventually approach (but do not reach) BH compactness – “deformable BSs are louder than stiff BHs”. Our conjecture for this result – supported by the numerical observations of “needles” in the GW power during mergers of unstable Branch IIUM mergers – is that during the collision phase, the increased deformation in a less compact BS results in a more strongly perturbed and asymmetric merger remnant whose subsequent relaxation generates more GWs. It will be interesting to see if this result extends to other types of horizonless compact objects such as Proca stars or perhaps even NSs.
@@
In general, BS models become unstable at high (but below BH) compactness where perturbations can trigger a rapid collapse to a BH. In both mini-BS and oscillaton models, this represents the only delineation point between stable and unstable BSs. However, in more complicated models with two or more minima in the potential, the presence of additional stable and unstable equilibrium points in the potential $V(|\phi_{\mathrm{c}}|)$ can introduce additional structure in the stability diagram (see e.g. Ref. Helfer:2016ljl). For the solitonic models we have studied, this manifests itself in the form of additional stable/unstable branches in the solution space of BSs (see Fig. 4). In these additional unstable regions, a BS may either collapse to a BH or migrate to a different BS. In all cases, the existence of these unstable regions implies that, in the limit of infinite infall time, the GW energy $E_{\rm GW}(|\phi_{\mathrm{c}}|)$ is discontinuous at the points of transition from stable to unstable branches. Unfortunately, since we do not measure GW energies over extraction spheres directly, but only the strain at one point on the sphere, there is scant hope for observing such discontinuities.
@@
In section IV, we have studied in detail the nature of the instability, in particular the time-scale or migration time of the transition. Furthermore, if the instability leads to a transition into another (excited) BS which oscillates, then we should expect “needles” structure in the energy function $E_{\rm GW}(|\phi_{\mathrm{c}}|)$. The main lesson here is that, there can be rapid changes in GW energy over small changes in the parameter space. For example, if we merge two equal mass oscillating BSs, the GW emitted would depend strongly on the phase of the pulsation at merger, complicating any observational inference over its progenitors. This raises two further important questions. (i) What is the timescale of relaxation of dynamic BS configurations to an equilibrium state? Numerical studies often result in long-lived dynamical stars (see e.g. Refs. Croft:2022bxq; Siemonsen:2023hko or our Fig. 5), but is it long enough to be of relevance for inspirals? (ii) What kind of environmental effects could induce dynamic perturbations of the kind experienced by our unstable migrating BSs?
One important point we would like to emphasise is that this work is focused on the mergers of in-phase BSs. However, as first investigated in Palenzuela:2007dm, it is well known that off-phase BSs exhibit significant deviation in the merger dynamics from in-phase BS. Broadly speaking, BSs with different phases exhibit repulsive forces with the magnitude depending on how off-phase they are Evstafyeva:2022bpr – this is due to the fact that BSs are wave-like objects and can interfere either constructively or destructively depending on their phases. It is unclear at the moment how these interference effects will affect mergers of unstable or near-unstable BSs. As a guess to one of its possible effects in BSs, we note that in off-phase mergers of real scalar Oscillatons Widdicombe:2019woy, it was shown that while off-phase repulsion resists the formation of BHs from direct mergers, the subsequent “compression” of individual Oscillatons due to the repulsion can cause sufficiently large disruption such that it collapses into a BH on its own. Thus, applying this logic to each BS, we anticipate that this might lead to a widening of the “red dots band” say in Fig. 1, although clearly there is plenty of scope for further investigation in this direction. Also, examining how waveforms change with the self-interaction parameter would provide a more complete perspective, and we leave this for future work.
@@
Finally, any program to search for BSs in the observational data stream requires the construction of waveform models or templates akin to the IMRPhenom, SEOBNR or NRSur approximants for BH-binary mergers. Our results suggest that the GWs from the BS star merger phase can exhibit significant differences in the GW amplitude as compared to BH mergers; cf. Fig. 9. The GW amplitude, however, is essentially degenerate with the luminosity distance and thus of little help for characterizing the nature of the source. This naturally raises the question whether the GW signature of the needle regime may exhibit more structural features in inspirals and/or for other BS parameters like unequal mass ratios or non-zero spins. Depending on the longevity of BS perturbations, comprehensive binary BS templates may need to account for these effects, perhaps using phenomenological parameters to extend simple semi-analytic quadrupole models (e.g. Ref. Toubiana:2020lzd) with rigid mass distributions. Given the complexity of the merger physics, however, we suspect that surrogate models, as for example constructed for Proca-star head-on mergers in Ref Luna:2024kof, may provide the most efficient way forward.
@@ Acknowledgements.
Figure 11: The $L_{2}$ norm of the Hamiltonian constraint as a function of time obtained for the BS binary configuration $\sigma_{0}=0.25$, $|\phi_{\rm c}|=0.02$ using different resolutions $N=64$, $N=256$ and $N=1024$.
Appendix A Constraint violations
The constraint violations observed in our simulations follow the pattern exemplified for the Hamiltonian constraint $H$ of the BS collision with parameters $\sigma=0.25$, $|\phi_{\rm c}|=0.02$ in Fig. 11. At early times, the constraint damping of the CCZ4 formulation reduces the constraint violations by an order of magnitude independent of numerical resolution. During the remainder of the infall, up to about t=300Mt=300\,M, the constraint violations of the $N=256$ and $N=512$ simulations nearly overlap. We attribute this to the fact that the constraint violations have very low magnitude and are initial-data dominated. To test this hypothesis, we have added a third simulation with lower resolution $N=64$. This indeed exhibits a significant contribution to $H$ arising from the discretization error with convergence of about second order. During merger, the constraint violations increase substantially, as expected for the formation of the extreme curvature region of a black hole  [^9] Note that this effect is enhanced by the effectively two-dimensional domain, which for purely geometric reasons attaches relatively higher weight to grid points near the black hole as compared to a three-dimensional grid.. Post merger, we observe convergence between first and second order.
Appendix B Gravitational Wave Extraction
Numerical relativity simulations in $3+1$ dimensions commonly use the Newman-Penrose scalar $\Psi_{4}$ to extract gravitational radiation. Here we follow the discussion of GW extraction in $d+1$ dimensions of Ref. Cook:2016qnt and apply it to our case of $2+1$ evolutions ($d=2$). The outgoing gravitational radiation is encoded in the the projections of the Weyl tensor $C_{ABCD}$ onto a null tetrad $\{k^{A},l^{A},m^{A},\bar{m}^{A}\}$
@@ Appendix B Gravitational Wave Extraction
here capital Latin indices run from 0 to $d$. In order to construct the null tetrad, we first define an orthonormal basis $\{e_{(0)}^{A},e_{(1)}^{A},e_{(2)}^{A},e_{(3)}^{A}\}$, where $e_{(0)}^{A}$ is the unit normal to the spatial hypersurfaces, $e_{(1)}^{A}$ is the unit normal radial vector and $e_{(2)}^{A},e_{(3)}^{A}$ are the angular vectors, which are typically constructed via Gram-Schmidt orthonormalisation. We then pick the tetrad consisting of ingoing and outgoing null vectors, $k^{A}$ and $l^{A}$ respectively, and a complex null vector $m^{A}$ and its conjugate $\bar{m}^{A}$,
@@
Following the notation of Ref. Cook:2016qnt, we will denote the off-domain components with a ’$w$’ index, so that in $2+1$ we have, $e_{(3)}^{A}\rightarrow e_{(w)}^{A}$. It is then straightforward to show that the outgoing gravitational radiation in Eq. (5) can be re-expressed by the real quantity
@@
where $\Omega_{22}$ and $\Omega_{ww}$ can be obtained directly from Eqs.(4.22)-(4.35) of Ref. Cook:2016qnt.